write an equation for the following function in vertex form:
then it shows a parabola,
the vertex is (2,1) and the point is (-2,5)

Answer by stanbon(75887)

write an equation for the following function in vertex form:
then it shows a parabola,
the vertex is (2,1) and the point is (-2,5)
------------------------------------------------
Form (x-h)^2 = 4p(y-k)
--------------------------
(h,k) is the vertex (2,1)
substitute to get:
(x-2)^2 = 4p(y-1)
------------------------
Let y = 5 when x = -2
(-2-2)^2 = 4p(5-1)
16 = 4p(4)
p = 1
----------------
Equation:
(x-2)^2 = 4(y-1)
========================
Cheers,
Stan H.